The length of a rectangle is 5 less than twice the width. The area of the rectangle is 7 square units. Find the dimensions of the rectangle. Hint: quadratic formula!
my work so far: 2w-5=L w(2w-5)=7

w(2w)-5w-7=0

Answer by checkley71(8403)

L=2W-5
W(2W-5)=7
2W^2-5W-7=0
(2W-7)(W+1)=0
2W-7=0
2W=7
W=7/2
W=3.5 ANSWR FOR THE WIDTH.
L=2*3.5-5
L=7-5
L=2
PROOF:
3.5(2*3.5-5)=7
3.5(7-5)=7
3.5*2=7
7=7